OpenStudy (anonymous):

i did it but it said incorrect any help thanks Use the quadratic formula to solve the following equation. 2a-5=4/a

OpenStudy (mrnood):

first multiply both sides of the equation by a write your result in your next answer...

OpenStudy (anonymous):

1/4 (5+/-r57)

OpenStudy (jdoe0001):

\(\bf 2a-5=\cfrac{4}{a}\implies 2a^2-5a=4\implies {\color{blue}{ 2}}a^2{\color{red}{ -5}}a{\color{green}{ -4}}=0 \\ \quad \\ \textit{quadratic formula}\\ \qquad \qquad x= \cfrac{ - {\color{red}{ b}} \pm \sqrt { {\color{red}{ b}}^2 -4{\color{blue}{ a}}{\color{green}{ c}}}}{2{\color{blue}{ a}}}\)

OpenStudy (anonymous):

2a-5=4/a 2a^2-5a-4=0

OpenStudy (anonymous):

so it is not a=14(5+57−−√),14(5−57−−√)

OpenStudy (jdoe0001):

plug in the values... what would they give you?

OpenStudy (mrnood):

@wildfire95 looks correct to me

OpenStudy (mrnood):

1/4 (5+/-r57) your first equation looks right (assuming r57 means sqrt(57)

OpenStudy (anonymous):

\[a=\frac{ 5\pm \sqrt{\left( -5 \right)^2-4*2*-4} }{ 2*2 }=\frac{ 5\pm \sqrt{25+32} }{4 }\] \[=\frac{ 5\pm \sqrt{57} }{ 4 }\]

OpenStudy (jdoe0001):

\(\bf 2a-5=\cfrac{4}{a}\implies 2a^2-5a=4\implies {\color{blue}{ 2}}a^2{\color{red}{ -5}}a{\color{green}{ -4}}=0 \\ \quad \\ \textit{quadratic formula}\\ \qquad \qquad x= \cfrac{ - {\color{red}{ -5}} \pm \sqrt { {\color{red}{ -5}}^2 -4{\color{blue}{ (2)}}{\color{green}{ (-4)}}}}{2{\color{blue}{ (2)}}}\implies \cfrac{5\pm\sqrt{57}}{4}\)